# Some insights into depth estimators for location and scatter in the multivariate setting

### v1

@@ 5 Appendix
Let $g\left(\delta\right)=\left(1-\delta\right)h\left(\frac{1}{3}\left(y_{\delta}+2\sqrt{y_{\delta}^{2}+6\ln 2}\right),y_{\delta}\right)$ with $y_{\delta}=\Phi^{-1}\left(\frac{\delta}{1-\delta}\right)$ a normal quantile, with $\delta<1/3.$Let $\varepsilon_{0}$ be the fixed point of $g$; i.e.,
$\varepsilon_{0}=g(\varepsilon_{0}).$ (6)
Lemma 3
The following statements hold.
  • (i)
  • (ii)
Proof. (i) Fixed $y_{0},$ we want to maximize $h_{0}\left(x\right)=h\left(x,y_{0}\right)=\Phi\left(x\right)-\Phi\left(\frac{y_{0}+x}{2}\right).$ If $\varphi$ stands for the derivative of $\Phi,$ the derivative of $h_{0}$turns out to be $h_{0}^{\prime}\left(x\right)=\varphi\left(x\right)-\frac{1}{2}\varphi\left(\frac{y_{0}+x}{2}\right)\geq 0$ if and only if $2exp\left(-\frac{1}{2}x^{2}\right)\geq exp\left(-\frac{1}{8}\left(\left(x+y_{0}\right)^{2}\right)\right)$ if and only if $\ln 2-\frac{3x^{2}}{8}+\frac{xy_{0}}{4}+\frac{y_{0}^{2}}{8}\geq 0$ if and only if $\frac{1}{3}\left(y_{0}-2\sqrt{y_{0}^{2}+6\ln 2}\right)$leq $x$leq $\frac{1}{3}\left(y_{0}+2\sqrt{y_{0}^{2}+6\ln 2}\right)$. Then $M\left(y_{0}\right)=\frac{1}{3}\left(y_{0}+2\sqrt{y_{0}^{2}+6\ln 2}\right)$ yields the maximum value given by
(ii) If $\varepsilon<\varepsilon_{0}$, using (6), we have that
$\varepsilon<(1-\varepsilon)h\left(M\left(y_{\varepsilon_{0}}\right),y_{\varepsilon_{0}}\right)=\frac{1-\varepsilon}{1-\varepsilon_{0}}(1-\varepsilon_{0}))h\left(M\left(y_{\varepsilon_{0}}\right),y_{\varepsilon_{0}}\right)=\frac{1-\varepsilon}{1-\varepsilon_{0}}\varepsilon_{0}.$
and the converse follows.
Lemma 4
If $y>-\sqrt{\ln\left(4\right),}$then $a>0,$ $a<a+b,$ $\varphi\left(a+b\right)<\varphi\left(a\right)$ and $f^{\prime}\left(y\right)<0,$ which entails the result. Since the function $y_{\delta}=\Phi^{-1}\left(\frac{\delta}{1-\delta}\right)$ is increasing in $\delta$ then the composition $f\left(y_{\delta}\right)=h\left(\frac{1}{3}\left(y_{\delta}+2\sqrt{y_{\delta}^{2}+6\ln 2}\right),y_{\delta}\right)$ is decreasing in $\delta$. Consequently, $g\left(y\right)$ is a product of two positive decreasing functions and it is decresing in $\delta$. Therefore, the claim of the Lemma follows.
Lemma 5
The following statements are valid.
  • (i)
    If $\varepsilon<1/5$ then there exists $-\infty<x_{L}<x_{U}<\infty$ such that
  • (ii)
    If there exists $0<\alpha\leq 1$ verifying that $\underset{x>0}{inf}$ $\underset{n\rightarrow\infty}{\underline{lim}}H_{n}\left([x,\infty\right))\geq\alpha$ and $\frac{\Phi\left(-\sqrt{\ln 4}\right)}{1+\Phi\left(-\sqrt{\ln 4}\right)}<\varepsilon<\varepsilon_{0},$ then there exists $-\infty<x_{L}<x_{U}<\infty$ such that
@@
  • (iii)
    If $P_{\varepsilon}\left(A^{(j)}\left(\hat{\mu}_{n},\hat{\sigma}_{n}\right)\right)\underset{n\rightarrow\infty}{\rightarrow}0$ for some $j\in\left\{1,2,3,4\right\}$ then $\underset{n\rightarrow\infty}{\overline{\lim}}D_{n}\left(\hat{\mu}_{n},\hat{\sigma}_{n}\right)$leq $\varepsilon$ and $\varepsilon>1/5.$
  • (iv)
    Given a sequence $\left\{\left(\mu_{n},\sigma_{n}\right)\right\}_{n\in\mathbb{N}}$ such that $\underset{n\rightarrow\infty}{\underline{\lim}}D_{n}\left(\mu_{n},\sigma_{n}\right)>0$ then $\underset{n\rightarrow\infty}{\underline{\lim}}P_{n}\left(A^{(j)}\left(\mu_{n},\sigma_{n}\right)\right)>0$ for any $j\in\left\{1,2,3,4\right\}.$
(ii) Given $y_{0}=y_{\varepsilon_{0}},$ and using Lemma 3 we can write $\ (1-\epsilon)h_{0}(x)$ $=$ $P_{\varepsilon}\left(A^{(3)}\left(\frac{y_{\varepsilon_{0}}+x}{2},\frac{x-y_{\varepsilon_{0}}}{2}\right)\right)$.The maximum of $h_{0}$ occurs at
and the maximum value is given by
If $\varepsilon<\varepsilon_{0}$, by (ii) of Lemma 3
Moreover
and then, $P_{n}\left(A^{(j)}\left(\frac{y_{\varepsilon_{0}}+M\left(y_{\varepsilon_{0}}\right)}{2},\frac{M\left(y_{\varepsilon_{0}}\right)-y_{\varepsilon_{0}}}{2}\right)\right)>\varepsilon$, $j=1,3$. Observe that $\frac{2}{3}y_{\delta}+\frac{1}{3}\sqrt{y_{\delta}^{2}+6\ln 2}>0$ if and only if $y_{\delta}\in\left[-\ln\left(4\right),\ln\left(4\right)\right]$ if and only if $\delta\geq\frac{\Phi\left(-\sqrt{\ln 4}\right)}{1+\Phi\left(-\sqrt{\ln 4}\right)}.$ Therefore $\frac{2}{3}y_{\varepsilon_{0}}+\frac{1}{3}\sqrt{y_{\varepsilon_{0}}^{2}+6\ln 2}>0,$and we can say that
then we get, for $n$ large enough that
and (ii) follows.
(iii) Take $-x_{L}$ obtained in item (i) and $j_{0}$ such that $\lim_{n}P_{\varepsilon}\left(A^{(j_{0})}\left(\hat{\mu}_{n},\hat{\sigma}_{n}\right)\right)=0.$ Then,
and the result follows.
(iv) It follows straightforwardly.
Lemma 6
If $\underset{x>0}{inf}$ $\underset{n\rightarrow\infty}{\underline{lim}}H_{n}\left([x,\infty\right))=1$ then there exists $-\infty<x_{L}<x_{U}<\infty$ such that
if and only if $\varepsilon<\varepsilon_{0}.$
Proof. If $\varepsilon<\varepsilon_{0}$ then Lemma (5) (i) and (ii) says the conclusion. If $\underset{x>0}{inf}$ $\underset{n\rightarrow\infty}{\underline{lim}}H_{n}\left([x,\infty\right))=1$ and there exists $-\infty<x_{L}<x_{U}<\infty$ such that $\varepsilon$ $<$  $\underset{n\rightarrow\infty}{\underline{\lim}}D_{n}\left(\frac{x_{L}+x_{U}}{2},\frac{x_{U}-x_{L}}{2}\right)$ then
$=$ $(1-\epsilon)h(x_{L},x_{U})$
$=$ $g(\epsilon).$
By Lemma 4 we get that $\varepsilon<\varepsilon_{0}.$
Lemma 7
Then,
(i) If $\lim_{n\rightarrow\infty}\mu_{n}-\sigma_{n}=-\infty$ then $\lim_{n\rightarrow\infty}P_{n}\left(A^{(1)}(\mu_{n},\sigma_{n})\right)=0$ and Lemma 5 (iv) yields a contradiction.
(ii) If $\mu_{n}+\sigma_{n}>a_{n}$ for $a_{n}$ large enough, then $\lim_{a\rightarrow\infty}P_{n}(A^{(1)}(\mu_{n},\sigma_{n}))=0$ and Lemma 5 (iv) yields a contradiction. Consequently,
$-\infty<\inf_{n}\left(\mu_{n}-\sigma_{n}\right)\text{ and }\mu_{n}+\sigma_{n}\leq a_{n}\text{ for }n\text{ sufficiently large.}$.
(iii) If $\mu_{n}+\sigma_{n}<a_{n}$ for $n$ sufficiently large then $\lim_{a\rightarrow\infty}P_{n}\left(A^{(3)}(\mu_{n},\sigma_{n})\right)=0,$ and Lemma 5 (iv) yields a contradiction. Then $\mu_{n}+\sigma_{n}=a_{n}$ for $n$ sufficiently large.
(iv) Summing up, we can assume without loss of generality that
By Lemma 6 there exists $-\infty<x_{L}<x_{U}<\infty$ such that
if and only if $\varepsilon<\varepsilon_{0}.$ Consequently, in case of having of point mass contaminations, $\left\{\left(\mu_{n},\sigma_{n}\right)\right\}_{n\in\mathbb{N}}$ cannot be the deepest estimators if and only if $\varepsilon<\varepsilon_{0}.$
Theorem 8
The breakdown point of the depth estimator is $\varepsilon^{\ast}=\varepsilon_{0}.$
Proof. Given $\varepsilon>0,$
for all $P\in\mathcal{F}_{\varepsilon}\left(\Phi\right)$. An special case is given by the point mass contaminations. Thus Lemma 7  let us infer that $\varepsilon<\varepsilon_{0}$ and the breakdown point verifies that $\varepsilon^{\ast}\leq\varepsilon_{0}.$
(ii) Assume that there is a sequence of distributions $P_{n}\in\mathcal{F}_{\varepsilon}\left(\Phi\right)$ such that the associated deepest estimators $\left\{\left(\hat{\mu}_{n}\hat{\sigma}_{n}\right)\right\}$ verifies that $\lim_{n\rightarrow\infty}\left(\hat{\mu}_{n}+\hat{\sigma}_{n}\right)=\infty.$ Set $A_{n}^{(j)}=A^{(j)}(\hat{\mu}_{n},\hat{\sigma}_{n}).$ Then $\lim_{n\rightarrow\infty}P_{\varepsilon}\left(A_{n}^{(j)}\right)=0$ for $j=4$ and Lemma 5 (iii) says that $\varepsilon>1/5.$ Without loss of generality we can assume that there is a -$\infty\leq b$ such that
Then
Suppose by contradiction that $\varepsilon<\varepsilon_{0}.$ By Lemma 5 (ii), there exists $-\infty<x_{L}<x_{U}<\infty$ such that
$\underset{n\rightarrow\infty}{\underline{lim}}D_{n}\left(\frac{x_{L}+x_{U}}{2},\frac{x_{U}-x_{L}}{2}\right)>\varepsilon\alpha$ (10)
Then, by (9) and (10), for $n$ large enough we have
which is a contradiction. Therefore $\varepsilon\geq\varepsilon_{0}$ and $\varepsilon^{\ast}\geq\varepsilon_{0}$ and the result follows.
Lemma 9
Given $\delta\in(0,1/2)$ let $\alpha=1-2\delta$ and $\varepsilon\in[0,1/2)$. Let $\mathcal{P}_{\varepsilon}\left(P_{0}\right)$ be the $\varepsilon-$contamination neighborhood with $P_{0}=N(\boldsymbol{\theta},\Sigma)$. Take $\mathcal{F}\left(M\right)$ as the set of symmetric and definite positive matrices $\Sigma$ such that the largest eigenvalue $\lambda_{1}(\Sigma)$ is less than a constant $M>0.$ For $\varepsilon\in\left[0,\varepsilon^{\prime}\right],$ $\varepsilon^{\prime}<1/3$ and $\left(p+\log\left(1/\delta\right)\right)/n$ sufficiently small, there exists a constant $C>0$ (depending on $\varepsilon^{\prime}$ but independent of $p,n,\varepsilon),$ such that, with probability greater than $1-2\delta,$ it holds that
$\inf_{\boldsymbol{\theta},\Sigma\in\mathcal{F}\left(M\right),P\in\mathcal{P}_{\varepsilon}\left(F_{0}\right)}P\left(\left\|\hat{\boldsymbol{\theta}}_{T}-\boldsymbol{\theta}\right\|^{2}\leq\tilde{C}\left(\max\left\{\frac{p}{n},B_{T}^{2}\left(\varepsilon,I\right)\right\}+\frac{\log\left(1/\delta\right)}{n}\right)\right)\geq\alpha.$
Proof. Chen et al. (2018) in the proof of Theorem 2.1 showed that, for $\mathbf{\theta}=\mathbf{0}$
$\Phi\left(\left\|\hat{\theta}\right\|\right)$$\leq$ $\frac{1}{2}+\frac{\varepsilon}{1-\varepsilon}+40\sqrt{\frac{6e\pi}{1-e^{-1}}}\sqrt{\frac{p+1}{n}}+\frac{7}{2}\sqrt{\frac{\log\left(1/\delta\right)}{n}}$
@@
$\left\|\hat{\theta}\right\|$$\leq$ $\Phi^{-1}\left(\frac{1}{2}+\frac{\varepsilon}{1-\varepsilon}+b\left(p,n\right)\right)-\Phi^{-1}\left(\frac{1}{2}\right)+\Phi^{-1}\left(\frac{1}{2}\right)$
@@
with $\text{with }\gamma_{C}$$\in$ $\left(\frac{1}{2},\frac{1}{2}+\frac{\varepsilon}{1-\varepsilon}+b\left(p,n\right)\right).$
Therefore, set and we get
$\left\|\hat{\theta}\right\|$$\leq$ $\Phi^{-1}\left(a\left(\varepsilon\right)+b\left(p,n\right)\right)-\Phi^{-1}\left(a\left(\varepsilon\right)\right)+\Phi^{-1}\left(a\left(\varepsilon\right)\right)$
@@
If $\varepsilon<1/3-c,$ $c$ a positive constant, $a\left(\varepsilon\right)$ is increasing on $[0,1/3-c)$, let $a(m_{c})$ be the maximum value. Take $d$ such that $01-a\left(m_{c}\right)>d>$ 0 and $\left(p,n\right)$in $A_{c,d}$ $=$ $\left\{\left(p,n\right):b\left(p,n\right)<1-a\left(m_{c}\right)-d\right\}.$Then, there exists a constant $C_{c,d}=\sup_{\left(p,n\right)\in A_{c,d}}$ $\left[\phi\left(\Phi^{-1}\left(a\left(m_{c}\right)+b\left(p,n\right)\right)\right)\right]^{-1}$ such that
$\left\|\hat{\theta}\right\|\leq\tilde{C}_{c,d}\left(\sqrt{\frac{p}{n}}\vee B_{L}\left(\mathbf{\hat{\theta}}_{T},\varepsilon,I\right)+\sqrt{\frac{\log\left(1/\delta\right)}{n}}\right),$
and the lemma follows.
Lemma 10
Given $\delta\in(0,1/2)$ let $\alpha=1-2\delta$ and $\varepsilon\in[0,1/2)$. Let $\mathcal{P}_{\varepsilon}\left(P_{0}\right)$ be the $\varepsilon-$contamination neighborhood with $P_{0}=N(\boldsymbol{\theta},\Sigma)$. Take $\mathcal{F}\left(M\right)$ as the set of symmetric and definite positive matrices $\Sigma$ such that the largest eigenvalue $\lambda_{1}(\Sigma)$ is less than a constant $M>0.$ For $\varepsilon\leq 1/3-c$ and $\left(p+\log\left(1/\delta\right)\right)/n$ sufficiently small, there exists a constant $C>0$ (depending on $\varepsilon^{\prime}$ but independent of $p,n,\varepsilon),$ such that, with probability greater than $1-2\delta,$
$\inf_{\Sigma\in\mathcal{F}\left(M\right),P\in\mathcal{P}_{\varepsilon}\left(F_{0}\right)}P\left(\left\|\hat{\Sigma}-\Sigma\right\|_{op}^{2}\leq C^{\ast}\left(\max\left\{\frac{p}{n},B_{E}^{2}(\varepsilon)\right\}+\frac{\log\left(1/\delta\right)}{n}\right)\right)\geq\alpha.$
Proof. In the process of obtaining (5), In the process of obtaining such an inequality, Chen et al. (2018, p. 1955) showed that, with probability greater than $1-2\delta,$
@@
If $\inf_{\mathbf{u\in}S^{p-1}}x_{\mathbf{u}}\leq\sup_{\mathbf{u}\in S^{p-1}}x_{\mathbf{u}}\leq 1$, $\sup_{\mathbf{u}\in S^{p-1}}\left|\Phi(\sqrt{\beta})-\Phi\left(\sqrt{\beta}x_{\mathbf{u}}\right)\right|=\Phi(\sqrt{\beta})-\Phi\left(\sqrt{\beta}\inf_{\mathbf{u}\in S^{p-1}}x_{\mathbf{u}}\right).$If $1\leq\inf_{\mathbf{u\in}S^{p-1}}x_{\mathbf{u}}\leq\sup_{\mathbf{u}\in S^{p-1}}x_{\mathbf{u}}$ then $\sup_{\mathbf{u}\in S^{p-1}}\left|\Phi(\sqrt{\beta})-\Phi\left(\sqrt{\beta}x_{\mathbf{u}}\right)\right|=\Phi\left(\sqrt{\beta}\sup_{\mathbf{u}\in S^{p-1}}x_{\mathbf{u}}\right)-\Phi(\sqrt{\beta}).$ If $\inf_{\mathbf{u\in}S^{p-1}}x_{\mathbf{u}}\leq 1\leq\sup_{\mathbf{u}\in S^{p-1}}x_{\mathbf{u}}$ we have to analyze both cases. If the maximum on the righthand side of (14) occurs in $\Phi\left(\sqrt{\beta}\sup_{\mathbf{u}\in S^{p-1}}x_{\mathbf{u}}\right)-\Phi(\sqrt{\beta})$ then $\sup_{\mathbf{u}\in S^{p-1}}x_{\mathbf{u}}\geq 1,$ and we have,
@@
00 0 $\leq$ $\sup_{u\in S^{p-1}}x_{u}-1\leq\frac{1}{2\sqrt{\beta}}\frac{1}{\phi\left(\Phi^{-1}\left(a\left(\varepsilon\right)+\frac{1}{2}a_{n,\delta}\right)\right)}a_{n,\delta}+\left[\frac{1}{\sqrt{\beta}}\Phi^{-1}\left(a\left(\varepsilon\right)\right)-1\right]=g_{S}\left(\varepsilon,n,\delta,p\right)$
@@
$B_{E}\left(\varepsilon\right)=\left[\frac{1}{\sqrt{\beta}}\Phi^{-1}\left(a\left(\varepsilon\right)\right)-1\right],\text{ }$
@@
00 0 $\leq$ $\Phi(\sqrt{\beta})-\Phi\left(\sqrt{\beta}\inf_{u\in S^{p-1}}x_{u}\right)\leq\frac{\varepsilon}{2\left(1-\varepsilon\right)}+\frac{1}{2}a_{n,\delta}$

### v2

@@ 7 Appendix
Proof of Lemma 3.1. Chen, Gao and Ren (2018a), in the proof of Theorem 2.1, showed that, for $\mathbf{\theta}=\mathbf{0}$
@@
$\left\|\hat{\theta}\right\| \displaystyle\leq$ $\Phi^{-1}\left(\frac{1}{2}+\frac{\varepsilon}{1-\varepsilon}+b\left(p,n\right)\right)-\Phi^{-1}\left(\frac{1}{2}\right)+\Phi^{-1}\left(\frac{1}{2}\right)$
@@
$\text{with }\gamma_{C}$ $\in$ $\left(\frac{1}{2},\frac{1}{2}+\frac{\varepsilon}{1-\varepsilon}+b\left(p,n\right)\right).$
Therefore, consider $\Phi^{-1}\left(a\left(\varepsilon\right)\right)$ rather than $\Phi^{-1}\left(\frac{1}{2}\right)$ to center $\Phi^{-1}\left(\frac{1}{2}+\frac{\varepsilon}{1-\varepsilon}+b\left(p,n\right)\right)$ and we get
$\left\|\hat{\theta}\right\| \displaystyle\leq$ $\Phi^{-1}\left(a\left(\varepsilon\right)+b\left(p,n\right)\right)-\Phi^{-1}\left(a\left(\varepsilon\right)\right)+\Phi^{-1}\left(a\left(\varepsilon\right)\right)$
@@
If $\varepsilon<1/3-c$, where $c$ is a positive constant, $a\left(\varepsilon\right)$ is increasing on [0,1/3c)[0,1/3-c), let $a(m_{c})$ be the maximum value. Take $d$ such that $1-a\left(m_{c}\right)>d>0$ and $\left(p,n\right)$ $\in$ $A_{c,d}$ == $\left\{\left(p,n\right):b\left(p,n\right)<1-a\left(m_{c}\right)-d\right\}.$ Consequently, there exists a constant
$C_{c,d}=\sup_{\left(p,n\right)\in A_{c,d}}\left[\phi\left(\Phi^{-1}\left(a\left(m_{c}\right)+b\left(p,n\right)\right)\right)\right]^{-1}$
such that
$\left\|\hat{\theta}\right\|\leq\tilde{C}_{c,d}\left(\sqrt{\frac{p}{n}}\vee B_{L}\left(\mathbf{\hat{\theta}}_{T},\varepsilon,\Phi\right)+\sqrt{\frac{\log\left(1/\delta\right)}{n}}\right),$
[…]
Proof of Lemma 3.2. In the process of obtaining (11), Chen, Gao and Ren (2018a) (p. 1955) showed that, with probability greater than $1-2\delta,$
@@
If $\inf_{\mathbf{u\in}S^{p-1}}x_{\mathbf{u}}\leq\sup_{\mathbf{u}\in S^{p-1}}x_{\mathbf{u}}\leq 1$, then
$\sup_{\mathbf{u}\in S^{p-1}}\left|\Phi(\sqrt{\beta})-\Phi\left(\sqrt{\beta}x_{\mathbf{u}}\right)\right|=\Phi(\sqrt{\beta})-\Phi\left(\sqrt{\beta}\inf_{\mathbf{u}\in S^{p-1}}x_{\mathbf{u}}\right).$
If $1\leq\inf_{\mathbf{u\in}S^{p-1}}x_{\mathbf{u}}\leq\sup_{\mathbf{u}\in S^{p-1}}x_{\mathbf{u}},$ then
$\sup_{\mathbf{u}\in S^{p-1}}\left|\Phi(\sqrt{\beta})-\Phi\left(\sqrt{\beta}x_{\mathbf{u}}\right)\right|=\Phi\left(\sqrt{\beta}\sup_{\mathbf{u}\in S^{p-1}}x_{\mathbf{u}}\right)-\Phi(\sqrt{\beta}).$
If $\inf_{\mathbf{u\in}S^{p-1}}x_{\mathbf{u}}\leq 1\leq\sup_{\mathbf{u}\in S^{p-1}}x_{\mathbf{u}}$ we have to analyze both cases. If the maximum on the right-hand side of (27) occurs in $\Phi\left(\sqrt{\beta}\sup_{\mathbf{u}\in S^{p-1}}x_{\mathbf{u}}\right)-\Phi(\sqrt{\beta})$ then $\sup_{\mathbf{u}\in S^{p-1}}x_{\mathbf{u}}\geq 1,$ and we have,
@@
$0$ $\leq$ $\sup_{u\in S^{p-1}}x_{u}-1$
$\leq$ $\frac{1}{2\sqrt{\beta}}\frac{1}{\phi\left(\Phi^{-1}\left(a\left(\varepsilon\right)+\frac{1}{2}a_{n,\delta}\right)\right)}a_{n,\delta}+\left[\frac{1}{\sqrt{\beta}}\Phi^{-1}\left(a\left(\varepsilon\right)\right)-1\right]=g_{S}\left(\varepsilon,n,\delta,p\right).$
@@
$B_{E}\left(\varepsilon\right)=\left[\frac{1}{\sqrt{\beta}}\Phi^{-1}\left(a\left(\varepsilon\right)\right)-1\right].\text{ }$
@@
$0$ $\leq$ $\Phi(\sqrt{\beta})-\Phi\left(\sqrt{\beta}\inf_{u\in S^{p-1}}x_{u}\right)\leq\frac{\varepsilon}{2\left(1-\varepsilon\right)}+\frac{1}{2}a_{n,\delta}$